The sum of two numbers is 12. Four times the smaller number is two less than the larger. What are the two numbers?
Answer by hamsanash1981@gmail.com(151)

let larger number be x and smaller number be y, then x + Y = 12 -- 1
and
4y = x-2 => x = 4y + 2
substituting x in equation 1
4y + 2 + y = 12
5y = 10
y = 2
when y = 2 then x = 12 - 2 => x = 10
the two numbers are 10 and 2.